OpenStudy (anonymous):

How do I solve problems like this? I have a math exam coming up, and I want to be prepared for solving problems like this. Please explain the process to me. Thanks! What are the exact solutions of x^2 - 5x - 7 = 0? I'll post the answer choices in the options. Thanks again!

OpenStudy (anonymous):

@ganeshie8 @mathmale @mathstudent55 @zepdrix @triciaal

ganeshie8 (ganeshie8):

familiar with quadratic formula ?

OpenStudy (anonymous):

A. \[x = \frac{ -5 \pm \sqrt3 }{ 2 }\] B. \[x = \frac{ 5 \pm \sqrt3 }{ 2 }\] C. \[x = \frac{ -5 \pm \sqrt53 }{ 2 }\] D. \[x = \frac{ 5 \pm \sqrt53 }{ 2 }\]

OpenStudy (anonymous):

@ganeshie8 I am

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] @ganeshie8 That's the quadratic equation, right?

OpenStudy (anonymous):

*formula

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (anonymous):

Oh, do I need to fill in the formula to solve the problem?

ganeshie8 (ganeshie8):

Yes

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

x^2 - 5x - 7 = 0 a = ? b = ? c = ?

OpenStudy (anonymous):

@ganeshie8 Sorry, for the late response, I'm helping someone else on their homework, I'll be back in about 1 minute, I'm almost done.

OpenStudy (anonymous):

@ganeshie8 Well that took 5 minutes, but I'm back

OpenStudy (anonymous):

a = 1 b = -5 c = -7

ganeshie8 (ganeshie8):

Looks good. Plug them in

OpenStudy (anonymous):

\[x = \frac{ -1 \pm \sqrt{1^2 - 4(1)(-7)} }{ 2(1) }\]

OpenStudy (anonymous):

1^2 = 1. -4*1 = -4. -4 * -7 = 28. 2 * 1 = 2

ganeshie8 (ganeshie8):

thats wrong, check once..

OpenStudy (anonymous):

\[x = \frac{ -1 \pm \sqrt{1 - 28} }{ 2 }\]

OpenStudy (anonymous):

1 - 28 = 27

ganeshie8 (ganeshie8):

b = -5 right ?

OpenStudy (anonymous):

Oh, that's wrong?

OpenStudy (anonymous):

Gaah OS IS FREAKING OUT. What did I get wrong @ganeshie8 ?

OpenStudy (anonymous):

Yeah, b = -5

OpenStudy (anonymous):

OHHHH

OpenStudy (anonymous):

\[x = \frac{ 5 \pm \sqrt{(-5)^2 - 4(1)(-7)} }{ 2() }\]

OpenStudy (anonymous):

I didn't even type in that order. wtf is going on @ganeshie8 ?

OpenStudy (anonymous):

Wait, now it's correct order again. NVM

OpenStudy (anonymous):

-5^2 = 25. 4ac = 28. 25 - 28 = -3. \[x = \frac{ 5 \pm \sqrt{-3} }{ 2 }\]

OpenStudy (anonymous):

@ganeshie8 Is that correct?

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (anonymous):

@ganeshie8 If u've already typed something, I can't see it. OS Is lagging like crazy and idk why

ganeshie8 (ganeshie8):

\[x = \frac{ 5 \pm \sqrt{(-5)^2 - 4(1)(-7)} }{ 2(1) } = \frac{5\pm \sqrt{25+28} }{2} \]

OpenStudy (anonymous):

Why 25 + 28? its b^2 - 4ac doesnt the - sign stick with the 4ac making it -4acac?

OpenStudy (anonymous):

*-4ac

OpenStudy (anonymous):

@ganeshie8 i know this isnt on my own post, but when youre done here do you think you could help me on my post?

OpenStudy (anonymous):

@cadenblah Just send him a message

OpenStudy (anonymous):

@ganeshie8 Help? Why would it be 25 + 28?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ganeshie8 Wanna go on PeeraAnswers or something?

OpenStudy (anonymous):

@ganeshie8 It doesn't matter I think I get how to solve the problem. I'll just use a calculator or something. I know how to get to the answer, and that's what matters. Thanks again!!!

OpenStudy (anonymous):

Refer to the attachment.